# Closed-Form Interpretation of Neural Network Classifiers with Symbolic Regression Gradients

### v1

@@ 5.1 Training a Neural Network For Binary Classification
I implemented one neural network architecture for all experiments using the Python library tensorflow [41]. This neural network consists of two hidden layers with 1000 neurons each and ELU activation functions. I compared different activation functions and ELU activations yielded gradients least affected by computational artifacts. The final layer contains one single neuron without a sigmoid activation function, commonly used for binary classification. The weights and biases in the hidden layers are regularized with an $L^{2}$ penalty of $10^{-4}$. Further, I employ dropout regularization after each hidden layer with a chance of $20\%$. I train the neural networks using the Adam optimizer and use learning rate and early stop callbacks that reduce the learning rate by 50% or stop training if the loss stops decreasing. For this reason, it is enough to set the number of epochs large enough such that the early stopping is always triggered. The batch size in all cases is 100. I train on $80\%$ of the data set and use another 20% for validation. Since I am not interested in calculating any test accuracy here, the validation and test set are the same.
Data: Labelled data set $D=(X_{train},Y_{train})$
Input: Neural Network Hyperparameters
1 Initialize neural network classifier with sigmoid activation at output $F$
2 Train $F$ on $D$
$f$leftarrow remove sigmoid activation of output neuron of $F$ Output: Trained model $F$, latent model $f$
Algorithm 1 Train Neural Network for Binary Classification
Data: Labelled data set $D=(X_{train},Y_{train})$
Data: Unlabelled data set $D_{u}=(X_{u})$triangleright Artificial data not used for training
Input: Trained model $F$ and latent model $f$
Data: Selection threshold $\delta$
1 $\tilde{X}\leftarrow(X_{train},X_{u})$triangleright Add artificial unlabelled data
2 $\tilde{X}\leftarrow\tilde{X}\textbf{ where }F(\tilde{X})\in[\delta,1-\delta]$triangleright Select data close to decision boundary
3 $G_{f}\leftarrow[\nabla f(\mathbf{x})\text{ for }\mathbf{x}\text{ in }\tilde{X}]$triangleright Gradients of latent model wrt. input
4 $G_{f}\leftarrow[\textbf{if }\nabla f(\mathbf{x})\neq 0:\nabla f(\mathbf{x})/\lVert\nabla f(\mathbf{x})\rVert$
@@ 1 $\tilde{X}\leftarrow(X_{train},X_{u})$triangleright Add artificial unlabelled data
Output: $(\tilde{X},G_{f})$
Algorithm 2 Obtain Gradients of Latent Model
Data: Gradient data set $(\tilde{X},G_{f})$
Input: Symbolic Regression Hyperparameters
Data: Set of unary and binary operations.
1 initialize symbolic regression model $T$
2 evolve $T$ with (
3 aaa $G_{T}\leftarrow[\nabla T(\mathbf{x})\text{ for }\mathbf{x}\text{ in }\tilde{X}]$triangleright Gradients of symbolic model
4 aaa $G_{T}\leftarrow[\textbf{if }\nabla T(\mathbf{x})\neq 0:\nabla T(\mathbf{x})/\lVert\nabla T(\mathbf{x})\rVert$
@@
6 aaa to minimize MSE$(G_{f},G_{T})$
Output: Symbolic regression model $T$
Algorithm 3 Symbolic Search
5.2 Obtaining Gradients of Latent Model
The second step is described in algorithm 2, which is used to collect the gradient information, see fig. 4(a), for the symbolic regression step. It involves potentially adding additional unlabelled data points to the training data set $X$, either from available unlabelled data, sampling from the data manifold, or by perturbing training data. This data set is denoted $\tilde{X}$. I omitted increasing the training data set in my experiments, but in data-scarce domains, this could lead to a significant improvement in the symbolic regression results.
@@ 5.2 Obtaining Gradients of Latent Model
Afterward, I calculate the normalized gradients $G_{f}=\nabla f/\lVert\nabla f\rVert$ of the latent model $f$, with respect to the input variables, on the remaining data points. Finally, together with the inputs they get stored in the form of a labeled data set $(\tilde{X},G_{f})$ that is used as training data of the symbolic regression algorithm.
5.3 Symbolic Search
The third step of the interpretation algorithm uses symbolic regression, see fig. 3, to find human-readable equations describing the decision boundary $g$. The algorithm is summarized in algorithm 3. This step employs the backend of the Python symbolic regression module PySR [31] coded in Julia called SymbolicRegression.jl.
The interpretation framework proposed in this paper has requirements beyond what symbolic regression algorithms typically do. Symbolic regression models are represented by trees $T$ that are trained on a labeled data set $(X,Y)$ to reproduce an output $Y$. However, the task in this manuscript is to perform symbolic regression on the normalized gradient $G_{f}=\nabla f/\lVert\nabla f\rVert$ of a specific neuron within the neural network. Thus, the gradient set $G_{f}$ takes on the role of the data set label $Y$. Further, the symbolic regression tree must be differentiable in order to calculate the normalized gradients of the tree $G_{T}=\nabla T/\lVert\nabla T\rVert$ itself. Finally, the desired solution of the symbolic regression procedure is not finding a closed form expression of the normalized gradient $\nabla f/\lVert\nabla f\rVert$, but the function $f$ itself. An additional limitation of SymbolicRegression.jl is its limitation to one-dimensional regression targets in the standard form.
@@ 5.3 Symbolic Search
I choose the mean square error (MSE) loss function between latent model normalized gradient $G_{f}$ and normalized gradient of the symbolic tree $G_{t}$ as the most straight-forward choice of the objective function to ascertain that both $f$ and $T$ fall into the same equivalence class defined by eq.(6). This loss function is equivalent to the cosine loss (CSL) based on the scalar product of two vectors.
@@
The output of the symbolic regression algorithm is not just a single closed-form function, but a set of symbolic functions that balance complexity and accuracy along the Pareto front. More precisely, the result contains functions with low complexity and low accuracy, and more complex functions are added if the complexity can be justified by an increase in accuracy.
Table 1: Data sets
@@
Experiment 1 2 $x_{1}^{2}+2x_{2}^{2}=1$
Experiment 2 2 $x_{1}^{2}+3x_{1}x_{2}+2x_{2}^{2}=5$
Experiment 3 3 $x_{1}^{2}+\sin(x_{2}+x_{3})=1$
Experiment 4 3 $\exp(x_{1}-x_{2})-\pi x_{3}=1$
Experiment 5 3 $x_{1}\exp(-\frac{1}{2}(x_{2}^{2}+x_{3}^{2}))=\frac{1}{5}$
Experiment 6 6 $\frac{x_{1}x_{2}}{(x_{3}-x_{4})^{2}+(x_{5}-x_{6})^{2}}=1$
6 Experiments
@@ 6.1 Data
I apply the outlined interpretation procedure to several neural networks that are tasked with classifying data sets that are each grouped into two classes. The classes are separated through a domain boundary defined by a decision formula $g$. The interpretation framework attempts to recover these functions from the neural networks. I perform experiments on six different data sets outlined in Table 1. For each data set, I sample 10000 data points and categorize them into classes according to the decision formulas. In all cases, the data sets are created with multiplicative Gaussian noise of $1\%$.
Figure 5: The results of fitting a symbolic regression model $T$ to the normalized gradients of the neural network are presented along the Pareto front. The Pareto front collects several possible results with decreasing Mean Square Error (MSE) and increasing complexity. The closest match to the true underlying function is often found at the point of steepest change of the Pareto front.
6.2 Results
Applying the proposed interpretation framework to the six data sets yields sets of formulas along the Pareto front. All formulas are equally valid and trade complexity for accuracy. The Pareto fronts for the six experiments are collected in fig. 5.
@@ 6.2 Results
7 Conclusion
@@
In this paper, I introduce a comprehensive framework to interpret neural network classifiers that is equally capable of interpreting single neurons in hidden layers or latent spaces. The solution presents the learned concept as a closed-form function. The interpretation method is based on the idea of finding an intersection between the space of human-readable functions and all functions that contain the same information as a specific neuron in an artificial neural network. Through the interpretation procedure, it is possible to get rid of uninterpretable transformations that conceal the true function and focus on the simplest form of the underlying concept.
@@
I demonstrate the power of the interpretation method in section 6 at the example of six different data sets with two to six input variables each. In 4 cases the interpretation procedure recovers the exact functional form of the true decision function, while in 2 experiments the procedure finds a very good symbolic approximation. Comparing these results with other works is impossible since most related works only interpret neural network regressors. However, the most similar research articles manage to only solve problems based on one variable [23] or two variables variables [24].
@@
In the future, it might be worth exploring how well a neural network classifier + interpretation approach compares to directly applying symbolic classification methods. I suspect that a neural network is a superior learner and the artificial data generation in algorithm 2 enables a better symbolic performance. I also believe that the neural network can deal with possible data noise better such that the artificially created data for the symbolic regression steps could be essentially noise-free.
There is of course the limitation of not being able to learn functions of a single variable since the interpretation method’s goal is to reduce the concept learned by a neural network to a one-dimensional function. This also means one loses the numerical value of the threshold/bias of the output neuron. This is however not a problem, since solving a one-dimensional symbolic regression problem or fitting a threshold/bias is minor compared to the complexity of problems solved by the presented interpretation framework.
It remains a central open question: why does a neuron contain the concept when it evaluates the function far away from the decision boundary? In my intuition, there is only a very weak training signal induced by inputs that are already classified correctly.
When applying the interpretation algorithm to practical problems it might be useful to sort out equations from the Pareto Front that violate dimensional constraints. Similarly, dimensional analysis could also be used inside the symbolic regression algorithm itself.
Further, one might have some inductive bias in mind that is better represented through other function sets. There is no need to use symbolic regression trees to represent the solution, any suitable differentiable function space can do.

### v2

@@ 3.3.1 Training a Neural Network For Binary Classification
I implemented one neural network architecture for all experiments using the Python library tensorflow [44]. This neural network consists of two hidden layers with 1000 neurons each and ELU activation functions. I compared different activation functions and ELU activations yielded gradients least affected by computational artifacts. The final layer contains one single neuron together with a sigmoid activation function, commonly used for binary classification. The weights and biases in the hidden layers are regularized with an $L^{2}$ penalty of $10^{-4}$. Further, I employ dropout regularization after each hidden layer with a chance of $20\%$. I train the neural networks using the Adam optimizer and use learning rate decay callbacks that reduce the learning rate by 50% loss stops decreasing. Further early stop callbacks stop the training process when converged. For this reason, it is enough to set the number of epochs large enough such that the early stopping is always triggered. The batch size in all cases is 100. I train on $80\%$ of the data set and use another 20% for validation. Since I am not interested in calculating any test accuracy here, the validation and test set are the same.
1
1) Train Neural Network for Binary Classification
2
1:Data: Labelled data set $D=(X_{train},Y_{train})$
2:Input: Neural Network Hyperparameters
3:Initialize neural network classifier with sigmoid activation at output $F$
4:Train $F$ on $D$
5:$f$leftarrow remove sigmoid activation of output neuron of $F$
6:Output: Trained model $F$, latent model $f$
52) Obtain Gradients of Latent Model
1:Data: Labelled data set $D=(X_{train},Y_{train})$
2:Data: Unlabelled data set $D_{u}=(X_{u})$triangleright Artificial data not used for training
3:Input: Trained model $F$ and latent model $f$
34:Input: Selection threshold $\delta$
5:$\tilde{X}\leftarrow(X_{train},X_{u})$triangleright Add artificial unlabelled data
6:$\tilde{X}\leftarrow\tilde{X}\textbf{ where }F(\tilde{X})\in[\delta,1-\delta]$triangleright Select data close to decision boundary
7:$G_{f}\leftarrow[\nabla f(\mathbf{x})\text{ for }\mathbf{x}\text{ in }\tilde{X}]$triangleright Gradients of latent model wrt. input
@@
10:Output: $(\tilde{X},G_{f})$
63) Symbolic Search
1:Data: Gradient data set $(\tilde{X},G_{f})$
2:Input: Symbolic Search Hyperparameters
3:Input: Set of unary and binary operations.
4:initialize symbolic model $T$
5:evolve $T$ with (
6:aaa $G_{T}\leftarrow[\nabla T(\mathbf{x})\text{ for }\mathbf{x}\text{ in }\tilde{X}]$triangleright Gradients of symbolic model
7:aaa $G_{T}\leftarrow[\textbf{if }\nabla T(\mathbf{x})\neq 0:\nabla T(\mathbf{x})/\lVert\nabla T(\mathbf{x})\rVert$
@@
79:aaa to minimize MSE$(G_{f},G_{T})$
10:Output: Symbolic model $T$
Algorithm 1 Neural Network Interpretation Algorithm
3.3.2 Obtaining Gradients of Latent Model
The second step is described in algorithm 1.2, which is used to collect the gradient information, see fig. 2a, for the symbolic search step. It involves potentially adding additional unlabelled data points to the training data set $X$, either from available unlabelled data, sampling from the data manifold, or by perturbing training data. This data set is denoted $\tilde{X}$. I omitted increasing the training data set in my experiments, but in data-scarce domains, this could lead to a significant improvement in the symbolic search results.
[…]
Afterward, I calculate the normalized gradients $G_{f}=\nabla f/\lVert\nabla f\rVert$ of the latent model $f$, with respect to the input variables, on the remaining data points. Finally, together with the inputs they get stored in the form of a labeled data set $(\tilde{X},G_{f})$ that is used as training data of the symbolic search algorithm.
3.3.3 Symbolic Search
The third step of the interpretation algorithm builds upon symbolic regression, see fig. 1c, to find human-readable equations describing the decision boundary $g$. The algorithm is summarized in algorithm 1.3. This step employs the backend of the Python symbolic regression module PySR [32] coded in Julia called SymbolicRegression.jl.
The interpretation framework proposed in this paper has requirements beyond what symbolic regression algorithms typically do. Symbolic regression models are represented by trees $T$ that are trained on a labeled data set $(X,Y)$ to reproduce an output $Y$. However, the task in this manuscript is to perform symbolic regression on the normalized gradient $G_{f}=\nabla f/\lVert\nabla f\rVert$ of a specific neuron within the neural network. Thus, the gradient set $G_{f}$ takes on the role of the data set label $Y$. Further, the symbolic tree must be differentiable in order to calculate the normalized gradients of the tree $G_{T}=\nabla T/\lVert\nabla T\rVert$ itself. Finally, the desired solution of the symbolic search procedure is not finding a closed form expression of the normalized gradient $\nabla f/\lVert\nabla f\rVert$, but the function $f$ itself. An additional limitation of SymbolicRegression.jl is its limitation to one-dimensional regression targets in the standard form.
@@ 3.3.3 Symbolic Search
I choose the mean square error (MSE) loss function between latent model normalized gradient $G_{f}$ and normalized gradient of the symbolic tree $G_{t}$ as the most straight-forward choice of the objective function to ascertain that both $f$ and $T$ fall into the same equivalence class defined by eq.4. This loss function is equivalent to the cosine loss (CSL) based on the scalar product of two vectors.
@@
The output of the symbolic search algorithm is not just a single closed-form function, but a set of symbolic functions that balance complexity and accuracy along the Pareto front. More precisely, the result contains functions with low complexity and low accuracy, and more complex functions are added if the complexity can be justified by an increase in accuracy.
4 Results
Table 1: Data sets are created by randomly sampling data points from multidimensional uniform distributions that are divided into two classes to represent binary classification problems.
@@ 4 Results
Experiment 1 2 $x_{1}^{2}+2x_{2}^{2}>1$
Experiment 2 2 $x_{1}^{2}+3x_{1}x_{2}+2x_{2}^{2}>5$
Experiment 3 3 $x_{1}^{2}+\sin(x_{2}+x_{3})>1$
Experiment 4 3 $\exp(x_{1}-x_{2})-\pi x_{3}>1$
Experiment 5 3 $x_{1}\exp(-\frac{1}{2}(x_{2}^{2}+x_{3}^{2}))>\frac{1}{5}$
Experiment 6 6 $\frac{x_{1}x_{2}}{(x_{3}-x_{4})^{2}+(x_{5}-x_{6})^{2}}>1$
Experiment 7 4 $x_{1}x_{2}>1\land x_{3}x_{4}>1$ vs $x_{1}x_{2}<1\land x_{3}x_{4}<1$
@@ 4.1 Data
I apply the outlined interpretation procedure to several neural networks that are tasked with classifying data sets that are each grouped into two classes. The classes are separated through a domain boundary defined by a decision formula $g$ and each data point is obtained by uniformly sampling within ranges that contain the decision boundary. The interpretation framework attempts to recover these functions from the neural networks. I perform experiments on different data sets outlined in table 1. For each data set, I sample 10000 data points and categorize them into classes according to the decision formulas. In all cases, the data sets are created with multiplicative Gaussian noise of $1\%$.
Figure 3: The results of fitting a symbolic model $T$ to the normalized gradients of the neural network are presented along the Pareto front. The Pareto front collects several possible results with decreasing Mean Square Error (MSE) and increasing complexity. The closest match to the true underlying function is often found at the point of steepest change of the Pareto front.
4.2 Experiments 1-6, Interpreting Decision Function
Applying the proposed interpretation framework to the data sets yields sets of formulas along the Pareto front. All formulas are equally valid and trade complexity for accuracy. The Pareto fronts for the experiments are collected in fig. 3.
@@ 4.2 Experiments 1-6, Interpreting Decision Function
These experiments can be compared to direct symbolic classification in cases where there is no ambiguity in what high-level feature is responsible for the class boundary as in experiments 1-6. Symbolic classification is based on symbolic regression together with a loss function that enables a categorical output. The application of symbolic classification and the corresponding results can be found in appendix A. It recovers three out of the six decision formulas together with almost exact values for the biases/thresholds.
4.3 Experiment 7, Symbolic Classification is not Neural Network Interpretation
Figure 4: a: Results of fitting a symbolic classification model to the data of experiment 7. b: Interpretation of a neural network classifying the same data set. c: Empirical correlation between symbolic classification, the proposed interpretation method, and latent model $f$. Symbolic classification learns a different high-level feature than the neural network. The interpretation framework presented in this paper correctly interprets the neural network.
Experiment 7 is different in the sense that the features are correlated and there are several possible concepts a model could learn to successfully classify the data. These correlations are common in machine learning data sets, eg when distinguishing between humans and dogs, one possible concept might be the presence of tails while another would be the color of the fur/skin. I train a neural network and interpret it similarly to the previous experiments. Further, I train a symbolic classification model to classify the data. In fig. 4a one can see that the symbolic classification model selects one of the high-level features $x_{1}x_{2}$ to make its decision. The interpretation procedure reveals that the neural network bases its prediction on a combination of the two high-level features $x_{1}x_{2}+x_{3}x_{4}$, see fig. 4b. In fig. 4c it is shown that the neural network has a very low correlation with either of the single features and confirms a very strong correlation between the concept learned by the neural network and the result of the interpretation framework. In this section, one can see that symbolic classification does not necessarily provide an interpretation of the neural network, whereas the framework presented in this paper does.
5 Discussion
@@
In this paper, I introduce a comprehensive framework for interpreting neural network classifiers that is equally capable of interpreting single neurons in hidden layers or latent spaces. The solution presents the learned concept as a closed-form function. The interpretation method is based on the idea of finding an intersection between the space of human-readable functions and all functions that contain the same information as a specific neuron in an artificial neural network. Through the interpretation procedure, it is possible to get rid of uninterpretable transformations that conceal the true function and focus on the simplest form of the underlying concept.
@@
I demonstrate the power of the interpretation method in section 4 at the example of different data sets with two to six input variables each. In 4 cases the interpretation procedure recovers the exact functional form of the true decision function, while in 2 experiments the procedure finds a very good symbolic approximation. In one experiment the neural network has the freedom of choosing between two different high-level features to make its prediction. The interpretation method reveals that the neural network learns a combination of both features.
Comparing these results with other works is impossible since most related works only replicate neural network regressors. As demonstrated in experiment 7, if there are multiple high-level concepts that can be employed to solve a specific task, replication is not interpretation. The most similar research articles manage to explain neural network regressors involving one variable [23] or two variables [24]. It is possible to compare the neural network interpretation technique to symbolic classification, see appendix A. Direct symbolic classification only finds three out of six closed-form expressions of the decision boundaries whereas gradient-based interpretation manages to find suitable solutions for all cases. There are multiple potential reasons for this performance difference. Neural networks are superior learners than genetic search algorithms. Symbolic search in the proposed interpretation framework is tasked with finding a closed-form expression extending deep into the full training manifold and hence experiences training signals from all data points. Direct symbolic classification only accumulates error signals from violations of the decision boundary. Neural networks are well suited to deal with noisy data such that the artificially created data of normalized gradients for the symbolic search steps could be essentially noise-free. Further, unlimited additional artificial data generation in algorithm 1.2 would further enable better symbolic performance.
@@
There is of course the limitation of not being able to learn functions of a single variable since the interpretation method’s goal is to reduce the concept learned by a neural network to a one-dimensional function. This also means one loses the numerical value of the threshold/bias of the output neuron. This is however not a problem, since solving a one-dimensional symbolic regression problem or fitting a threshold/bias is a minor challenge compared to the complexity of problems solved by the presented interpretation framework.
When applying the interpretation algorithm to practical problems it might be useful to sort out equations from the Pareto Front that violate dimensional constraints. Similarly, dimensional analysis could also be used inside the symbolic search algorithm itself. Further, one might have some inductive bias in mind that is better represented through other function sets. There is no need to use symbolic trees to represent the solution, any suitable differentiable function space can do. Lastly, this interpretation method is more useful in bottleneck layers and when neurons are disentangled, since the framework cannot capture information distributed among multiple neurons.
This article is accompanied by the article ”Closed-Form Interpretation of Neural Network Latent Spaces with Symbolic Gradients” [45], where we show how to interpret latent spaces of neural networks in closed form.
@@ 6 Acknowledgements
Appendix A Symbolic Classification
Figure 5: The results of fitting a symbolic classification model $T$ to six experiments. The Pareto front collects several possible results with decreasing Mean Square Error (MSE) and increasing complexity. The closest match to the true underlying function is often found at the point of steepest change of the Pareto front.
The most closely related technique to learn what neural network classifiers learn is symbolic classification. However, it is important to note that symbolic classification is not an interpretation technique, it can only provide equivalent results if there is only one possible high-level concept that a neural network could learn in order to make its predictions. Symbolic classification is based on symbolic regression with hinge/margin loss function, more precisely, the loss function I use is
$\mathcal{L}(F(X),Y)=\frac{1}{n}\sum_{i=1}^{n}\max(0,1-(y_{i}F(\mathbf{x}_{i})))$ (21)
I perform symbolic classification on experiments 1-6 with the same symbolic regression algorithm and hyperparameters as the experiments in the main body of this paper. The model is trained using the following building blocks: input variables $x_{1},...,x_{n}$, floating-point constants, binary operators: +, -, *, /, unary operators: sin, exp. The hyperparameters are a batch size of 25, a maximum number of epochs of 200, an early stop callback, and a maximum tree size of 30 nodes.
The Pareto fronts describing the results can be found in fig. 5. In three experiments 1, 3, 4 the symbolic classification algorithm finds the exact decision boundary, while in experiments 2, 5, 6 the algorithm fails to discover anything meaningful. Symbolic classification has the advantage of also recovering the threshold/bias of the last layer in contrast to the interpretation framework presented in this manuscript. However, symbolic regression adds increased complexity to its algorithm because it spends nodes on the exact determination of the bias/threshold. Further, the gradient-based interpretation framework has the freedom to choose from a larger function set and can thus provide more simplifications in its solutions.